INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM FLUORIDE IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, KIMBERLY HANDLER, RACHEL A. HARRIS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, MEREDITH REDICK, BECCA STEINBERG, Department of Chemistry and Biochemistry and Center for Nanoscience, University of Missouri, St Louis, MO 63121-4499.

Several new bands of PtF has been recorded in the near-infrared with rotational resolution using intracavity laser absorption spectroscopy. The results of the analysis will be presented. The gas phase PtF molecules were produced using a platinum-lined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.